OpenStudy (anonymous):

Find the indefinite integral of 1/(24x^27)

OpenStudy (amistre64):

\[\int\frac{1}{ax^n}dx;n\ne1\] \[\frac{1}{a(n-1)x^{n-1}}\] if i recall it correctly

OpenStudy (anonymous):

i think we can do it this way = 1 * x^-27 -- 24 INT = 1 * x^-26 -- ----- + c 24 -26 (using INT a x^m dx = a x ^(m+1) -------- + c ) m + 1